OpenStudy (anonymous):

what is 1/2 * 1/sqrt{(2+sqrt5)/2} ?

OpenStudy (anonymous):

\[\frac{ 1 }{2 } *\frac{ 1 }{\sqrt{(2+\sqrt{5})/2 }}\]

OpenStudy (lgbasallote):

\[\huge \frac 12 \times \frac 1{\sqrt{\frac{(2 + \sqrt 5)}2}}\] 2 is under the root too?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@lgbasallote: help me plz